OpenStudy (akimi):

how do i subtract two powers... for example 4^8-8 both 8s are powers

OpenStudy (saadsword):

4^(8-8)= 4^(0)= 1

OpenStudy (akimi):

wow its that simple thanks i wasnt sure

OpenStudy (saadsword):

no problem

OpenStudy (akimi):

so if \[16\times2^{m}=4^{8-8}\] would m equal \[\sqrt{7.5}\]

jhonyy9 (jhonyy9):

what is there please 16 time 2^m = ?

OpenStudy (akimi):

im solving for M

OpenStudy (akimi):

the question was If n = 8 and 16 • 2m = 4n – 8, then m =

jhonyy9 (jhonyy9):

on the right side 4^8-8 = 4^0 = 1 so from this result 16*2^m = 1 ohhh than there are 16*2^m = 4^n -8 here in the right side just n is the exponent of 4 ?

jhonyy9 (jhonyy9):

ok. i understand that this is your question but from there just m is the power of 2 or n is exponent too ?

jhonyy9 (jhonyy9):

16 = 2^4 2^4 *2^m = 2^(4+m) this is for left side but what is exactly on the right side ?

jhonyy9 (jhonyy9):

on the right side n is exponent same with m on the left side ?

OpenStudy (akimi):

when you plug \[n=8\] in the right side \[4^{n-8}\] you get 1 after that they ask you to solve \[m\] which is \[16\times2^{m}\] it ends up being \[16\times2^{m}=1\]

OpenStudy (leonhard_euler):

\[4^{8-8} = 4^{8}\times4^{-8} = \frac{ 4^{8} }{ 4^{8} } = 1\]